Parth (parthkohli):

OwlKid's kiddy proof. Show that an integer to an even power is a perfect square.

Parth (parthkohli):

\( \color{Black}{\Rightarrow START}\) Let's start out with the identity \(\large (a^b)^c = a^{bc}\). Wait. Let's leave that hanging above for a while. A prerequisite for all people is to know the definition of even numbers. An even number is an integer in the form \(2k: k \in \mathbb{Z}\). Let's assume that we have \( x^{2k}\), where both \(k\) and \(x\) are integers. \( \color{Black}{\Rightarrow x^{2k} }\) Use that identity NOW. \( \color{Black}{\Rightarrow (x^k)^2}\) Now, if you sense, we have a perfect square. Remember - \((\cdots)^2\) is a perfect square! HENCE, PROVED. $HAKA!

hero (hero):

Question says "show" not "prove". Not really sure this is a "proof" :P

Parth (parthkohli):

Additional note: A perfect square is a number which can be expresses as \((\mathbb{Z})^2\), and in order for \((\cdots)^2\) to be a perfect one, \(\cdots\) must be a square.

Parth (parthkohli):

Well, it is one of them. Showing is just proving ;)

OpenStudy (uri):

NO ITS NOT @ParthKohli

Parth (parthkohli):

A debate here? I don't need it.

OpenStudy (uri):

K iw as kidding SHOWING MEANS PROVING HAPPY!

hero (hero):

@uri, you're so cute :P

OpenStudy (uri):

Well,i am! :D why does everyone say that anyway?

hero (hero):

Not everyone knows how be cute, so when they see someone who is, they feel compelled to acknowledge it :P

Parth (parthkohli):

:|

OpenStudy (uri):

K thanks @Hero :)

OpenStudy (zzr0ck3r):

lol how is that not a proof? I guess you should have started with the words assume or let then it would "look" more like a proof.

Parth (parthkohli):

That is why I am kiddy :)

OpenStudy (zzr0ck3r):

i guess you shoult state the universe but outside of that looks fine to me:)

Parth (parthkohli):

'Kiddy proofs' is like the own series that I started to state axioms. There is no proof. I agree. I sometimes pity on the people who ask 'why so?' to questions like these. There's where I decided to make this kind of stuff.

Parth (parthkohli):

I hope you understand what I feel :)